OpenStudy (anonymous):

Questions on geometric distribution Alec and Bill take alternate turns at kicking a football at a goal, and their probabilities of scoring a goal on each kick are p1 and p2 respectively, independently of previous outcomes. The first person to score allows the other person one more kick. If the other then scores, the game is drawn. If the other then misses, the first has won the game. Alec begins a game. Given that p1 = p2 = 1/3, find the expected total number of kicks in the game. [2] Answer is 4 but I had no idea why. Anyone care to explain? (with calculations possibly)

OpenStudy (anonymous):

Other info given is Show that the probability that (i) Alec scores first, on his nth kick, is p1(q1q2)^(n−1), where q1 = 1 − p1 and q2 = 1 − p2, [2] (ii) Alec wins the game is p1q2/ (1 − q1q2), [3] (iii) the game is drawn is p1p2(1 + q1)/(1 − q1q2) . [3]

OpenStudy (anonymous):

ok we can do this, lets go slow so i can remember how as we go. if i recall correctly you do not actually sum a geometric series, but rather the derivative of one

OpenStudy (anonymous):

FYI, I have no prob showing the all 3 probabilities. But I couldn't for the rest of my life figure out the last question. Any help would be appreciated. :)

OpenStudy (anonymous):

oh ok so we do not have show the probabilities from scratch, just compute the expectation. i should write this on a piece of paper before we start, first thing we do is replace the p by 1/3 and q by 2/3 so we can see what we actually have

OpenStudy (anonymous):

i) \[\frac{1}{3}(\frac{2}{9})^{n-1}\] ii)\[\frac{2}{5}\] iii) \[\frac{1}{3}\] is what i get, is that what you have?

OpenStudy (anonymous):

got the same thing.

OpenStudy (anonymous):

i am trying to figure out how to use them but i am starting from scratch. how can the game end in 2 kicks? goal, not 1/3*2/3 goal goal 1/3*1/3 3 kicks not goal not (2/3)^2(1/3) not goal goal 2/3 (1/3)^2 4 kicks no no goal goal (1/3)^2(2/3)^2 no no goal no 1/3(2/3)^3 does this look right?

OpenStudy (anonymous):

ok i see a pattern emerging finally

OpenStudy (anonymous):

i think we have so sum two series, one where there is a win and one where there is a draw.

OpenStudy (anonymous):

win 2 kicks \[\frac{1}{3}\frac{2}{3}\] 3 kicks \[(\frac{2}{3})^2\frac{1}{3}\] 4 kicks \[(\frac{2}{3})^3\frac{1}{3}\] n kicks \[( \frac{2}{3}) ^{n-1}\frac{1}{3}\]

OpenStudy (anonymous):

now we have to multiply by the number of kicks and add this up \[\frac{1}{3}\left(2\frac{2}{3}+3\frac{2}{3}^2+4\frac{2}{3}^3+...\right)\]

OpenStudy (anonymous):

inside sum is 8 although that might not be obvious, so the answer so far is \[\frac{8}{3}\] hmm that means the other one must only by 1/4

OpenStudy (anonymous):

i dont see why would the 4 kicks differ from the 2 kicks...

OpenStudy (anonymous):

they are essentially the same thing.

OpenStudy (anonymous):

now for "draw" 2 kicks, \[\frac{1}{9}\] 3 kicks \[\frac{2}{3}\frac{1}{9}\] 4 kicks \[(\frac{2}{3})^2\frac{1}{9}\] n kicks \[(\frac{2}{3})^{n-2}\frac{1}{9}\]

OpenStudy (anonymous):

if i read this correctly there are two possibitlites for it ending with 4 kicks, one ending in a win and the other ending in a tie no no goal no no no goal goal

OpenStudy (anonymous):

first one has probability (2/3)^3(1/3) second on (2/3)^2(1/3)^2

OpenStudy (anonymous):

ok maybe i am being stupid and we can add these up

OpenStudy (anonymous):

second try probability it ends in 2 kicks is 1/3 3 kicks 2/3 4 kicks 4/27 5 kicks 8/81

OpenStudy (anonymous):

so looks like we are summing up \[2(\frac{1}{3})(\frac{2}{3})^0+3(\frac{1}{3})(\frac{2}{3})^1+4\frac{1}{3}(\frac{2}{3})^2+...)\] where the \[=\frac{1}{3}\left(2\times( \frac{2}{3})^0+3(\frac{2}{3})^1+(\frac{2}{3})^2+...\right)\]

OpenStudy (anonymous):

and sure enough, the answer is \[\frac{1}{3}\times 12=4\]

OpenStudy (anonymous):

i don't see why shouldn't we calculate the win for Bill... your 'win' is essentially only counting Alex right?

OpenStudy (anonymous):

bill can win too right?

OpenStudy (anonymous):

yes. and the draw.

OpenStudy (anonymous):

oh sorry, yes i was calculating the win for both bill wins on odd kicks

OpenStudy (anonymous):

for example bill can win on no goal no

OpenStudy (anonymous):

so there are two ways for there to be any number of kicks, win or draw odd kicks bill can win, even ones alex can win

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

i can write down what i computed if you like, i computed two columns, one for win and one for draw, then added them

OpenStudy (anonymous):

but what did you do with the 8/3? i don't see it appearing anywhere in the calculations below

OpenStudy (anonymous):

that all was a mistake, i was trying to add the series up separately, before i realized it would be easier to add them together

OpenStudy (anonymous):

i was going to add up one series for wins, and one for draws, but it was foolish, i didn't see the pattern when i added them

OpenStudy (anonymous):

ok got it.

OpenStudy (anonymous):

ok but it doesn't explain how to add \[\sum n(\frac{2}{3})^{n-2}\] though

OpenStudy (anonymous):

if you know how to do it, then we can take it as a given. it requires a trick actually two tricks

OpenStudy (anonymous):

pretty sophisticated question i would say best of the day so far

OpenStudy (anonymous):

maybe i will reword and and repost it as a challenge

OpenStudy (anonymous):

Sry, where did u get the 2(1/3)(2/3)^0+3(1/3)(2/3)+13/3(2/3)^2+...) from ?

OpenStudy (anonymous):

As for the sum you would probably need to substitute it with binomial expansion to calculate.

OpenStudy (anonymous):

oh ok lets go slow 2 kicks probabilty is 1/3*2/3+1/3*1/3= 1/3 3 kicks 2/3*1/3*2/3+2/3*1/3*1/3 =1/3*2/3 4 kicks 2/3*2/3*1/3*2/3 2/3*2/3*1/3*1/3 = 1/3*4/9 etc

OpenStudy (anonymous):

no wait but i added up the equations for n kicks and sum up the thing it doesnt adds to the same amount.

OpenStudy (anonymous):

i saw the pattern because the numbers were 1/3, 2/9,4/27,8/81 but in order to make a nice goemetric series i had to factor out a 1/3 to make it work

OpenStudy (anonymous):

as for summing \[\sum_{k=2}^{\infty}n(\frac{2}{3})^{n-2}\] you can use a nice trick

OpenStudy (anonymous):

instead of summing that one , find the following sum \[\sum_{n=1}^{\infty}n(\frac{2}{3})^{n-1}\] which looks exactly like the derivative of \[\sum_{n =0}^{\infty} (\frac{2}{3})^n\]

OpenStudy (anonymous):

problem is i don't get (2/3)^(n-2) by adding up (1/3)(2/3)^(n-1) and (2/3)^(n-2)(1/9)

OpenStudy (anonymous):

hold on you lost me there

OpenStudy (anonymous):

i just added and saw the pattern there are two ways to end in n kicks a win which has probability \[(\frac{2}{3})^{n-1}\frac{1}{3}\] and a loss which has probability \[(\frac{2}{3})^{n-2}\frac{1}{3}\frac{1}{3}\] when you add you get \[\frac{1}{3}(\frac{2}{3})^{n-2}(\frac{2}{3}+\frac{1}{3})\] \[\frac{1}{3}(\frac{2}{3})^{n-2}\]

OpenStudy (anonymous):

oh i see no you do not get \[(\frac{2}{3})^{n-2}\] you get \[\frac{1}{3}(\frac{2}{3})^{n-2}\]

OpenStudy (anonymous):

ok i got it apparently i did\[\sum_{n=1}^{\infty} n(3/4)(2/3)^n \] instead of \[\sum_{n=2}^{\infty} n(3/4)(2/3)^n\]

OpenStudy (anonymous):

No wonder i couldn't get the answer. Anyway, Thx for the help.

OpenStudy (anonymous):

not sure where the \[\frac{3}{4}\] comes from

OpenStudy (zarkon):

it is the same as yours sat

OpenStudy (anonymous):

oh pehaps another trick better than the one i used. expand

OpenStudy (anonymous):

i mean explain (please)

OpenStudy (zarkon):

\[\frac{1}{3}(\frac{2}{3})^{n-2}=\frac{1}{3}(\frac{2}{3})^{n}\frac{9}{4}=(\frac{2}{3})^{n}\frac{3}{4}\]

OpenStudy (anonymous):

oh yes, a much better trick i see

OpenStudy (anonymous):

oh no actually i don't see, i was summing \[\sum nx^{n-1}=\frac{1}{(1-x)^2}\] what for \[\sum nx^n\] ?

OpenStudy (anonymous):

oh i expanded them first then used the same trick 1/(1-x)^2. cept that i don't really recognize the summation form of the binomial expansion.

OpenStudy (anonymous):

ok good night then. its 2am in my time. thx for everything.

OpenStudy (zarkon):

\[\sum nx^n=x\sum nx^{n-1}\]

OpenStudy (anonymous):

pretty tough stuff for only 2 marks

OpenStudy (anonymous):

ok but then you are really after \[\sum nx^{n-1}\] so why not do that at the first step @Zarkon

OpenStudy (zarkon):

you can

OpenStudy (zarkon):

If this is in reference to a couple posts above..I was just showing where his 3/4 came from...not that I would do that

OpenStudy (anonymous):

ok i just don't know why you would multiply to get \[nx^n\] and then divide to find what you want. seems like an odd way to do business